# thanks: These authors contribute equallythanks: These authors contribute equally

### v1

@@ Abstract
Physical reservoir computing provides a powerful machine learning paradigm that exploits nonlinear physical dynamics for efficient information processing. By incorporating quantum effects, quantum reservoir computing gains superior potential in machine learning applications, for the quantum dynamics are exponentially costly to simulate classically. Here, we present a novel quantum reservoir computing approach based on correlated quantum spin systems, exploiting natural quantum many-body interactions to generate reservoir dynamics, thereby circumventing the practical challenges of deep quantum circuits. Our experimental implementation supports nontrivial quantum entanglement and exhibits sufficient dynamical complexity for high-performance machine learning. We achieve state-of-the-art performance in experiments on standard time-series benchmarks, reducing prediction error by one to two orders of magnitude compared to previous quantum reservoir experiments. In long-term weather forecasting, our 9-spin quantum reservoir delivers greater prediction accuracy than classical reservoirs with thousands of nodes. This represents a first experimental demonstration of quantum machine learning outperforming large-scale classical models on real-world tasks.
@@
Introduction.—The rapid growth of artificial intelligence has highlighted the demand for novel computational paradigms capable of real-time processing of temporally structured, data-intensive information streams. Physical reservoir computing (PRC) offers a promising solution by harnessing the intrinsic dynamics of physical substrates as computational resources within a machine learning framework Tanaka et al. (2019); Nakajima (2020). In this approach, inputs are projected into the high-dimensional space of a fixed dynamical system that simultaneously performs processing and memory, thus avoiding the von Neumann bottleneck. The output is obtained by only training a linear readout layer, enabling fast inference with minimal learning cost. PRC can be implemented in a broad range of physical systems that provide a high-dimensional state space, nonlinear input–output mapping, and fading memory Konkoli (2017), including memristors Moon et al. (2019); Milano et al. (2022); Liu et al. (2022); Liang et al. (2024), spintronics Torrejon et al. (2017); Romera et al. (2018); Grollier et al. (2020); Lee et al. (2024), and optical systems Brunner et al. (2013); Vandoorne et al. (2014); Larger et al. (2017); Rafayelyan et al. (2020). As larger and more complex dynamical systems tend to enhance computational capacity Kaplan et al. (2020); Raghu et al. (2017); Lukoševičius and Jaeger (2009), this framework naturally extends into the quantum regime. Quantum reservoir computing (QRC) leverages the exponentially large Hilbert space and complex correlations of quantum systems to offer unique advantages in solving challenging tasks. Given that some noisy intermediate-scale quantum (NISQ) systems have already outperformed classical simulators Arute et al. (2019); Zhong et al. (2020); Madsen et al. (2022), current quantum devices hold strong potential for realizing QRC with computational capacity and expressiveness beyond classical models.
@@
Despite its strong potential, QRC has so far been explored primarily through theoretical studies and numerical simulations on various idealized quantum models, while experimental realizations remain relatively limited Mujal et al. (2021); Ghosh et al. (2021). Most implementations rely on deep quantum circuits to generate the complex dynamics required for QRC Hu et al. (2024); Pfeffer et al. (2022); Chen et al. (2020); Kubota et al. (2023); Suzuki et al. (2022); Yasuda et al. (2023); Monzani et al. (2024). However, such circuits often demand substantial depth, making them vulnerable to gate errors and dissipation, thereby limiting the fidelity and performance of the QRC in the experiments. Furthermore, while quantum systems inherently possess exponentially large Hilbert spaces, current readout schemes are typically limited to local measurements, restricting the full exploitation of the high-dimensional nature of the quantum dynamics.
@@
In this Letter, we present an experimental realization of quantum reservoir computing based on a quantum spin system. To address the abovementioned challenges, we introduce two key innovations: the integration of both coherent spin-spin interactions and incoherent spin relaxation processes to enrich dynamical complexity of the reservoir, and the use of time-multiplexed measurements to increase the effective readout capacity of the reservoir Fujii and Nakajima (2017). Building on these, we achieve the state-of-the-art performance in QRC experiments on the widely used NARMA benchmark task, reducing the prediction error by one to two orders of magnitude compared to previous results Suzuki et al. (2022); Yasuda et al. (2023); Monzani et al. (2024). We apply the QRC protocol to weather forecasting, which is crucial for countless social and economic activities Price et al. (2025), yet challenging due to the chaotic nature of the atmosphere Kalnay (2003); Palmer (1993). In long-term predictions, our 9-spin quantum reservoir achieves higher prediction accuracy than classical reservoirs with thousands of nodes. Our results suggest practical quantum advantages in time-series prediction may be achievable with current quantum hardware. To our knowledge, this represents a first experimental demonstration of a quantum machine learning system that outperforms large-scale classical networks on realistic datasets, where previous advantages were typically shown via numerical simulations or limited to synthetic datasets Biamonte et al. (2017); Huang et al. (2021); Liu et al. (2021); Cerezo et al. (2022); Caro et al. (2022); Jäger and Krems (2023); Xia et al. (2023); Kornjača et al. (2024).
@@
Model design.—We introduce the general framework of QRC for temporal information processing, as illustrated in Fig. 1. Consider a pair of discrete scalar sequences: the input $\{s_{k}\}$ and the target output $\{y_{k}\}$, where $y_{k}$ depends on a finite input history up to time step $k$, denoted by $y_{k}=f(s_{k},\ldots,s_{k-\Delta})$. The goal of QRC is to approximate this relationship by generating an output sequence $\{\hat{y}_{k}\}$ that closely matches the target. The inputs $\{s_{k}\}$ are encoded into a control sequence and sequentially applied to an $N$-qubit quantum reservoir, mapping the information into a $4^{N}$-dimensional Hilbert space. Between successive inputs, the reservoir undergoes a fixed intrinsic evolution, which integrates new inputs with the memory of past inputs. To extract the information, specific observables are measured at each time step, yielding a sequence of readout vectors ${\bf x}_{k}$} with $x_{k}^{i}={\rm Tr}(\rho_{k}O^{i})$. The output is computed as a weighted sum of readout components, $\hat{y}_{k}={\bf x}_{k}^{\top}{\bf w}$, with a bias term incorporated by augmenting ${\bf x}_{k}$ with $x^{0}_{k}=1$. Following standard machine learning practices, the readout weights $\bf w$} are optimized on a training dataset by minimizing the error $\sum_{k=1}^{L}\left(y_{k}-\hat{y}_{k}\right)^{2}$, and the performance of the resulting QRC is then evaluated on test data.
@@
Figure 2: Schematic of experimental setup and readout schemes. (a) Circuit diagram of the QRC experimental implementation. The reservoir, composed of carbon and proton spins, undergoes repeated cycles of input-dependent rotations $\{\phi_{k},\theta_{k}\}$ followed by intrinsic evolution under $\exp(\tau\mathcal{L}_{H})$. The experiment adopts a rewinding protocol, where each experiment at time step $k$ is restarted from cycle $k-n_{\rm wo}$ to wash out initial state dependence, and readout is performed on the proton spins at the end of the circuit. Two readout schemes are considered: (b) Single-time readout: analogous to capturing multiple images of a static system by artificially changing the viewpoint. A fixed set of observables $\{O^{i}\}$ (e.g. Pauli operators) is measured on the reservoir state $\rho_{k}$, yielding the readout vector ${\bf x}_{k}=[\langle O^{1}\rangle_{k},\langle O^{2}\rangle_{k},\dots]^{\top}$ where $\langle O\rangle_{k}={\rm Tr}(\rho_{k}O)$. (c) Time-multiplexed readout: analogous to filming a moving system from a fixed viewpoint, where other aspects naturally rotate into view. A single observable $O$ is sampled at multiple time points during the reservoir’s evolution, equivalently measuring the Heisenberg‑picture operators $\{O(t_{i})\}$, yielding the readout vector ${\bf x}_{k}=[\langle O(t_{1})\rangle_{k},\langle O(t_{2})\rangle_{k},\dots]^{\top}$.
@@
The computational capabilities of QRC essentially depend on both the dynamics of the quantum reservoir and the number of readout components Mart´ınez-Peña et al. (2021); Xia et al. (2022); Dambre et al. (2012); Mart´ınez-Peña et al. (2023). Generally, QRC dynamics consist of unitary and non-unitary components. The unitary evolution, realized via analog Hamiltonian dynamics or digital quantum circuits, transforms and spreads information throughout the system while preserving coherence and generating entanglement. In addition to enriching the dynamics, the non-unitary part is essential for retaining memory of recent inputs by gradually diminishing the influence of past states, a property known as fading memory Boyd and Chua (1985). Previous protocols have achieved this through qubit resetting Fujii and Nakajima (2017); Hu et al. (2024) or engineered dissipative channels Chen et al. (2020); Sannia et al. (2024), but these methods often require intricate control techniques or additional experimental resources.
@@
Here, we adopt a more efficient method by leveraging the inherent relaxation, which commonly exists in natural quantum systems. Thus, the system evolution is governed by the Lindbladian superoperator $\dot{\rho}=\mathcal{L}_{H}[\rho]=-i\left[H,\rho\right]+\mathcal{R}[\rho]$, where $H$ is the system’s Hamiltonian and $\mathcal{R}$ denotes the relaxation channel. The overall state update is then expressed as
@@
where $e^{\tau\mathcal{L}_{H}}$ denotes the propagator for evolution over duration $\tau$, and $\mathcal{U}_{s_{k}}[\rho]=U_{s_{k}}\rho U^{\dagger}_{s_{k}}$ represents the unitary channel used to encode the classical input $s_{k}$. In spin systems, two dominant relaxation processes are longitudinal relaxation (with time constant $T_{1}$), which drives the system toward thermal equilibrium via energy dissipation, and transverse relaxation (with time constant $T_{2}$), which causes coherence decay through phase randomization Levitt (2008). These processes correspond to generalized amplitude damping and dephasing channels, respectively Nielsen and Chuang (2001). Our results indicate that $T_{1}$ relaxation plays a pivotal role in ensuring QRC performance (see Supplementary Material sup ), in agreement with recent numerical studies Domingo et al. (2023); Kubota et al. (2023); Monzani et al. (2024). By incorporating inevitable relaxation as a computational resource, our method overcomes coherence time limitations and eliminates the need for qubit resets, enabling prolonged experimental temporal processing with minimal control.
@@
Despite the rich dynamics of quantum reservoirs, the information processing capacity of a dynamical system is fundamentally limited by the number of independent readout functions Dambre et al. (2012). While quantum systems provide access to an exponentially large state space, extracting complete information through state tomography becomes increasingly challenging for large systems, particularly when high-order correlations are involved Nielsen and Chuang (2001). Consequently, most existing schemes focus on extracting local observable information like $\langle\sigma_{z}^{i}\rangle$ Chen et al. (2020); Suzuki et al. (2022); Kubota et al. (2023); Yasuda et al. (2023); Monzani et al. (2024), which capture only a small subset of the reservoir’s degrees of freedom, thus failing to fully exploit its computational potential.
To address this limitation, we employ time multiplexing, a technique that effectively increases the number of readouts by performing repeated measurements at regular intervals during the reservoir’s evolution Fujii and Nakajima (2017), as illustrated in Fig. 2(c). When input encoding and evolution are decoupled in QRC, time multiplexing can be viewed as a form of temporal tomography, where fixed observables are measured and information about other components is continuously mapped onto them through evolution Yang et al. (2020); Chantasri et al. (2020); Ghosh et al. (2020). In our implementation, we utilize the free induction decay signal, a natural manifestation of time multiplexing in quantum systems, to capture temporal information, with experimental results showing significant performance improvement.
@@
High-accuracy temporal prediction.—To demonstrate the performance of our approach, we perform experimental benchmarks on a nuclear magnetic resonance (NMR) quantum platform. The initial task we consider is the nonlinear autoregressive moving average (NARMA) simulation, a widely adopted benchmark for temporal processing capabilities Hochreiter (1997); Atiya and Parlos (2000). In accordance with previous work Fujii and Nakajima (2017); Suzuki et al. (2022); Yasuda et al. (2023); Monzani et al. (2024), the input sequence is constructed as a superposition of sine waves, while the target sequence is generated by specific recurrent dynamics, for example:
@@
where $n$ determines the number of past time steps influencing the current output, with the task denoted by NARMAnn. Here, we allocate 400 steps for training and 100 steps for testing. Readout weights are optimized via ridge regression Hoerl and Kennard (1970): ${\bf w}^{*}=\left({\bf X}^{\top}{\bf X}+\lambda{\bf I}\right)^{-1}{\bf X}^{\top}{\bf y}$, where ${\bf X}=\left[{\bf x}_{1},\ldots,{\bf x}_{400}\right]^{\top}$ and ${\bf y}=\left[y_{1},\ldots,y_{400}\right]^{\top}$ denote the readout matrix and target vector for training, and $\lambda$ is the regularization strength. By fitting separate $\bf w$$^{*}$ for each $n$ on the same $\bf X$}, the reservoir simultaneously emulates multiple NARMA systems, a capability commonly referred to as multitasking Fujii and Nakajima (2017).
Fig. 2(a) presents the circuit schematic for our QRC experiments. The quantum reservoir is a 9-spin coupling system consisting of four carbon nuclei and five proton nuclei in a 13C-labeled crotonic acid molecule. The spin reservoir undergoes successive cycles of input encoding and free evolution according to Eq. (1). For NARMA task, its scalar input is normalized to $\bar{s}_{k}\in[0,1]$ and encoded via an $x$-axis radio-frequency pulse on proton nuclei, effecting a global rotation $U_{s_{k}}=\prod_{n=5}^{9}R^{(n)}_{x}\left(\arcsin(\bar{s}_{k})\right)$. The system Hamiltonian is
@@
where $\nu_{i}$ represents the chemical shift of spin-$i$ and $J_{ij}$ denotes the coupling between spin-$i$ and spin-$j$ sup . To circumvent the measurement back-action, we adopt the rewinding protocol Mujal et al. (2023), where the experiment of time step $k$ is performed with input cycles from $k-n_{\rm wo}$ to $k$. The first $n_{\rm wo}$ cycles serve to wash out the initial state dependence under the premise of fading memory, thus allowing the experiment to start from any state without specific preparation. The washout length $n_{\rm wo}$ depends on the system’s relaxation rates and evolution durations, and is determined through preliminary experiments sup .
After the $n_{\rm wo}+1$ cycles, readout is performed on the proton spins using two schemes, as illustrated in Fig. 2(b) and (c). Single‑time readout refers to the conventional measurement of a fixed set of observables, here chosen as single-qubit Pauli operators. As $\rm H_{3}$$\rm H_{5}$ are indistinguishable methyl protons, only nine distinct expectation values, $\langle\sigma^{m}_{\alpha}\rangle$ ($\alpha=x,y,z$, $m={\rm H_{1}},{\rm H_{2}},{\rm methyl}$), are accessible. Time-multiplexed readout is realized using the free induction decay (FID) signal, which uniformly samples the nuclear transverse magnetization during free evolution Levitt (2008):
$S(t)\propto{\rm Tr}\left\{e^{t\mathcal{L}_{H}}\left[U\rho U^{\dagger}\right]\,\sum_{m=5}^{9}\left(\sigma^{m}_{y}+i\sigma^{m}_{x}\right)\right\}.$ (4)
In the experiments, we apply a $\pi/2$ pulse before FID readout ($U=\prod_{m=5}^{9}R^{m}_{x}(\pi/2)$) and apply a Fourier transform to the FID signal to isolate the reservoir response and reduce dimensionality, yielding 653 spectral features. Further details are provided in Supplementary Material sup .
Figure 3: High accuracy temporal prediction with NARMA benchmark. (a) Prediction performance (evaluated by $R^{2}$) on NARMA tasks with $n$ varying from 2 to 20, using quantum reservoirs with different configurations. Red and blue lines represent evolution durations of $\tau=$0.3\text{\,}\mathrm{s}$\$ and $0.01\text{\,}\mathrm{s}$, respectively, while solid and dashed lines indicate results obtained with time multiplexing (TM) and single-time readout schemes. (b) Prediction results for NARMA tasks with $n=10$ and $20$, utilizing time-multiplexed readout with $\tau=$0.01\text{\,}\mathrm{s}$\$.
@@
where $\bar{y}=\frac{1}{L}\sum_{k=1}^{L}y_{k}$. For direct comparison with prior work, we also report the corresponding normalized mean square errors in TABLE 1; our best cases achieve error reductions of 1-2 orders of magnitude relative to recent experimental benchmarks Suzuki et al. (2022); Yasuda et al. (2023); Monzani et al. (2024). Time multiplexing markedly improves performance over single-time readout, especially for $\tau=$0.01\text{\,}\mathrm{s}$\$, where predictions are nearly perfect across all NARMA orders (see NARMA10 and 20 in Fig. 3(b) and additional results in Supplementary Material sup ). As the relaxation effects become more pronounced with longer evolution times, the information from high-order correlations is significantly diminished at $\tau=$0.3\text{\,}\mathrm{s}$\$, thereby reducing the benefit of time multiplexing. In the Supplementary Material, it is shown that the quantum reservoir significantly outperformed its classical spin-based counterpart, which yields an average $R^{2}\textless 0.67$ across NARMA2 to NARMA20, highlighting the role of quantum correlations in computation sup .
@@
Realistic weather forecasting.—To show the practical utility of our approach, we conduct weather forecasting on a daily climate dataset from Delhi, a benchmark commonly employed in the machine learning community dai (2019). Our analysis focuses on predicting the multivariate time series of temperature and humidity, ${\bf y}_{k}=\left[{\rm Temp}(k),{\rm Humi}(k)\right]$. We first perform single-step-ahead forecasting, where ${\bf y}_{k}$ serves as the $k$-th input and ${\bf y}_{k+1}$ as the corresponding target. The dataset is divided into 374 washout steps, 600 training steps, and 600 testing steps. The normalized temperature and humidity data are encoded as angles $\theta_{k}$ and $\phi_{k}$ for global rotations applied to the proton and carbon nuclei, respectively, with an input interval of $\tau=$0.03\text{\,}\mathrm{s}$\$. Fig. 4(a) presents the single-step forecasting results alongside the target values from the test set. Real weather data exhibit substantial stochastic fluctuations, which pose challenges for precise forecasting. Nevertheless, QRC predictions successfully capture the overall trends and demonstrate close agreement with local variations in several segments of the time series.

### v2

@@ Abstract
Physical reservoir computing provides a powerful machine learning paradigm that exploits nonlinear physical dynamics for efficient information processing. By incorporating quantum effects, quantum reservoir computing offers superior potential for machine learning applications, as quantum dynamics are exponentially costly to simulate classically. Here, we present a novel quantum reservoir computing approach based on correlated quantum spin systems, exploiting natural quantum many-body interactions to generate reservoir dynamics, thereby circumventing the practical challenges of deep quantum circuits. Our experimental implementation supports nontrivial quantum entanglement and exhibits sufficient dynamical complexity for high-performance machine learning. We achieve state-of-the-art performance in experiments on standard time-series benchmarks, reducing prediction error by 1 to 2 orders of magnitude compared to previous quantum reservoir experiments. In long-term weather forecasting, our 9-spin quantum reservoir delivers greater prediction accuracy than classical reservoirs with thousands of nodes. This represents the first experimental demonstration of quantum machine learning outperforming large-scale classical models on real-world tasks.
@@
Introduction—The rapid growth of artificial intelligence has highlighted the demand for novel computational paradigms capable of real-time processing of temporally structured, data-intensive information streams. Physical reservoir computing (PRC) offers a promising solution by harnessing the intrinsic dynamics of physical substrates as computational resources within a machine learning framework [79, 60]. In this approach, inputs are projected into the high-dimensional space of a fixed dynamical system that simultaneously performs processing and memory, thus avoiding the von Neumann bottleneck. The output is obtained by only training a linear readout layer, enabling fast inference with minimal learning cost. PRC can be implemented in a broad range of physical systems that provide a high-dimensional state space, nonlinear input–output mapping, and fading memory [37], including memristors [56, 54, 47, 46], spintronics [80, 71, 22, 44], and optical systems [9, 82, 43, 67]. As larger and more complex dynamical systems tend to enhance computational capacity [32, 68, 49], this framework naturally extends into the quantum regime. Quantum reservoir computing (QRC) leverages the exponentially large Hilbert space and complex correlations of quantum systems to offer unique advantages in solving challenging tasks. Given that some noisy intermediate-scale quantum (NISQ) systems have already outperformed classical simulators [2, 88, 50], current quantum devices hold strong potential for realizing QRC with computational capacity and expressiveness beyond classical models.
@@
Despite its strong potential, QRC has so far been explored primarily through theoretical studies and numerical simulations on various idealized quantum models, while experimental realizations remain relatively limited [58, 20]. Most implementations rely on deep quantum circuits to generate the complex dynamics required for QRC [26, 65, 12, 41, 77, 86, 55]. However, such circuits often demand substantial depth, making them vulnerable to gate errors and dissipation, thereby limiting the fidelity and performance of the QRC in the experiments. Furthermore, while quantum systems inherently possess exponentially large Hilbert spaces, current readout schemes are typically limited to local measurements, restricting the full exploitation of the high-dimensional nature of the quantum dynamics.
@@
In this Letter, we present an experimental realization of quantum reservoir computing based on a quantum spin system. To address the abovementioned challenges, we introduce two key innovations: the integration of both coherent spin-spin interactions and incoherent spin relaxation processes to enrich the dynamical complexity of the reservoir, and the use of time-multiplexed measurements to increase the effective readout capacity of the reservoir [19]. Building on these, we achieve the state-of-the-art performance in QRC experiments on the widely used NARMA benchmark task, reducing the prediction error by 1 to 2 orders of magnitude compared to previous results [77, 86, 55]. We apply the QRC protocol to weather forecasting, which is crucial for countless social and economic activities [66], yet challenging due to the chaotic nature of the atmosphere [31, 62]. In long-term predictions, our 9-spin quantum reservoir achieves higher prediction accuracy than classical reservoirs with thousands of nodes. Our results suggest practical quantum advantages in time-series prediction may be achievable with current quantum hardware. To our knowledge, this represents a first experimental demonstration of a quantum machine learning system that outperforms large-scale classical networks on realistic datasets, where previous advantages were typically shown via numerical simulations or limited to synthetic datasets [5, 27, 48, 11, 10, 30, 83, 38].
@@
Model design—We introduce the general framework of QRC for temporal information processing, as illustrated in Fig. 1. Consider a pair of discrete scalar sequences: the input $\{s_{k}\}$ and the target output $\{y_{k}\}$, where $y_{k}$ depends on a finite input history up to time step $k$, denoted by $y_{k}=f(s_{k},\ldots,s_{k-\Delta})$. The goal of QRC is to approximate this relationship by generating an output sequence $\{\hat{y}_{k}\}$ that closely matches the target. The inputs $\{s_{k}\}$ are encoded into a control sequence and sequentially applied to an $N$-qubit quantum reservoir, mapping the information into a $4^{N}$-dimensional Hilbert space. Between successive inputs, the reservoir undergoes a fixed intrinsic evolution, which integrates new inputs with the memory of past inputs. To extract the information, specific observables are measured at each time step, yielding a sequence of readout vectors ${\bf x}_{k}$} with $x_{k}^{i}={\rm Tr}(\rho_{k}O^{i})$. The output is computed as a weighted sum of readout components, $\hat{y}_{k}={\bf x}_{k}^{\top}{\bf w}$, with a bias term incorporated by augmenting ${\bf x}_{k}$ with $x^{0}_{k}=1$. Following standard machine learning practices, the readout weights $\bf w$} are optimized on a training dataset by minimizing the error $\sum_{k=1}^{L}\left(y_{k}-\hat{y}_{k}\right)^{2}$, and the performance of the resulting QRC is then evaluated on test data.
@@
Figure 2: Schematic of experimental setup and readout schemes. (a) Circuit diagram of the QRC experimental implementation. The reservoir, composed of carbon and proton spins, undergoes repeated cycles of input-dependent rotations $\{\phi_{k},\theta_{k}\}$ followed by intrinsic evolution under $\exp(\tau\mathcal{L}_{H})$. The experiment adopts a rewinding protocol, where each experiment at time step $k$ is restarted from cycle $k-n_{\rm wo}$ to wash out initial state dependence, and readout is performed on the proton spins at the end of the circuit. Two readout schemes are considered. (b) Single-time readout: analogous to capturing multiple images of a static system by artificially changing the viewpoint. A fixed set of observables $\{O^{i}\}$ (e.g. Pauli operators) is measured on the reservoir state $\rho_{k}$, yielding the readout vector ${\bf x}_{k}=[\langle O^{1}\rangle_{k},\langle O^{2}\rangle_{k},\dots]^{\top}$ where $\langle O\rangle_{k}={\rm Tr}(\rho_{k}O)$. (c) Time-multiplexed readout: analogous to filming a moving system from a fixed viewpoint, where other aspects naturally rotate into view. A single observable $O$ is sampled at multiple time points during the reservoir’s evolution, equivalently measuring the Heisenberg‑picture operators $\{O(t_{i})\}$, yielding the readout vector ${\bf x}_{k}=[\langle O(t_{1})\rangle_{k},\langle O(t_{2})\rangle_{k},\dots]^{\top}$.
@@
The computational capabilities of QRC essentially depend on both the dynamics of the quantum reservoir and the number of readout components [51, 84, 16, 52]. Generally, QRC dynamics consist of unitary and nonunitary components. The unitary evolution, realized via analog Hamiltonian dynamics or digital quantum circuits, transforms and spreads information throughout the system while preserving coherence and generating entanglement. In addition to enriching the dynamics, the nonunitary part is essential for retaining memory of recent inputs by gradually diminishing the influence of past states, a property known as fading memory [6]. Previous protocols have achieved this through qubit resetting [19, 26] or engineered dissipative channels [12, 73], but these methods often require intricate control techniques or additional experimental resources.
@@
Here, we adopt a more efficient method by leveraging the inherent relaxation which commonly exists in natural quantum systems. Thus, the system evolution is governed by the Lindbladian superoperator $\dot{\rho}=\mathcal{L}_{H}[\rho]=-i\left[H,\rho\right]+\mathcal{R}[\rho]$, where $H$ is the system’s Hamiltonian and $\mathcal{R}$ denotes the relaxation channel. The overall state update is then expressed as
@@
where $e^{\tau\mathcal{L}_{H}}$ denotes the propagator for evolution over duration $\tau$, and $\mathcal{U}_{s_{k}}[\rho]=U_{s_{k}}\rho U^{\dagger}_{s_{k}}$ represents the unitary channel used to encode the classical input $s_{k}$. In spin systems, two dominant relaxation processes are longitudinal relaxation (with time constant $T_{1}$), which drives the system toward thermal equilibrium via energy dissipation, and transverse relaxation (with time constant $T_{2}$), which causes coherence decay through phase randomization [45]. These processes correspond to generalized amplitude damping and dephasing channels, respectively [61]. Our results indicate that $T_{1}$ relaxation plays a pivotal role in ensuring QRC performance (see Supplemental Material [1]), in agreement with recent numerical studies [17, 41, 55]. By incorporating inevitable relaxation as a computational resource, our method overcomes coherence time limitations and eliminates the need for qubit resets, enabling prolonged experimental temporal processing with minimal control.
@@
Despite the rich dynamics of quantum reservoirs, the information processing capacity of a dynamical system is fundamentally limited by the number of independent readout functions [16]. While quantum systems provide access to an exponentially large state space, extracting complete information through state tomography becomes increasingly challenging for large systems, particularly when high-order correlations are involved [61]. Consequently, most existing schemes rely on measuring local observables like $\langle\sigma_{z}^{i}\rangle$ [12, 77, 41, 86, 55], which capture only a small subset of the reservoir’s degrees of freedom, thus failing to fully exploit its computational potential.
To address this limitation, we employ time multiplexing, a technique that effectively increases the number of readouts by measuring fixed observables at different times $\{t_{i}\}$ during the reservoir’s evolution [19], as illustrated in Fig. 2(c). By accessing the Heisenberg-evolved observables $\{O(t_{i})\}$, this scheme provides more independent information about the reservoir state than a single-time measurement can offer, thereby raising the upper bound on the achievable performance [1]. In our implementation, we utilize the free induction decay (FID) signal, a natural time-multiplexed readout in quantum systems, with experimental results showing significant performance improvements.
@@
High-accuracy temporal prediction—To demonstrate the performance of our approach, we perform experimental benchmarks on a nuclear magnetic resonance (NMR) quantum platform. The initial task we consider is the nonlinear autoregressive moving average (NARMA) simulation, a widely adopted benchmark for temporal processing capabilities [23, 3]. In accordance with previous work [19, 77, 86, 55], the input sequence is constructed as a superposition of sine waves, while the target sequence is generated by specific recurrent dynamics, for example:
@@
where $n$ determines the number of past time steps influencing the current output, with the task denoted by NARMAnn. We allocate 400 steps for training and 100 steps for testing. Readout weights are optimized via ridge regression: ${\bf w}^{*}=\left({\bf X}^{\top}{\bf X}+\lambda{\bf I}\right)^{-1}{\bf X}^{\top}{\bf y}$, where ${\bf X}=\left[{\bf x}_{1},\ldots,{\bf x}_{400}\right]^{\top}$ and ${\bf y}=\left[y_{1},\ldots,y_{400}\right]^{\top}$ denote the readout matrix and target vector for training, and $\lambda$ is the regularization strength [24]. By fitting separate $\bf w$$^{*}$ for each $n$ on the same $\bf X$}, the reservoir simultaneously emulates multiple NARMA systems, a capability referred to as multitasking [19].
Figure 2(a) presents the circuit schematic for our experiments. The quantum reservoir is a 9-spin coupling system consisting of four carbon nuclei and five proton nuclei in a 13C-labeled crotonic acid molecule. The spin reservoir undergoes successive cycles of input encoding and free evolution according to Eq. (1). For NARMA task, its scalar input is normalized to $\bar{s}_{k}\in[0,1]$ and encoded via an $x$-axis radio-frequency pulse on proton nuclei, effecting a global rotation $U_{s_{k}}=\prod_{n=5}^{9}R^{n}_{x}\left[\arcsin(\bar{s}_{k})\right]$, where $R^{n}_{x}(\theta)=e^{-i\frac{\theta}{2}\sigma^{n}_{x}}$ and $\sigma^{n}_{x,y,z}$ are Pauli operators on spin $n$. The system Hamiltonian is
@@
where $\nu_{i}$ represents the chemical shift of spin $i$ and $J_{ij}$ denotes the coupling between spin $i$ and spin $j$ [1]. To circumvent the measurement backaction, we adopt the rewinding protocol [57], where the experiment of time step $k$ is performed with input cycles from $k-n_{\rm wo}$ to $k$. The first $n_{\rm wo}$ cycles serve to wash out the initial state dependence under the premise of fading memory, thus allowing the experiment to start from any state without specific preparation. The washout length $n_{\rm wo}$ depends on the system’s relaxation rates and evolution durations, and is determined through preliminary experiments [1].
After the $n_{\rm wo}+1$ cycles, readout is performed on the proton spins using two schemes, as illustrated in Figs. 2(b) and (c). Single‑time readout refers to the conventional measurement of a fixed set of observables, here chosen as single-qubit Pauli operators. As $\rm H_{3}$$\rm H_{5}$ are indistinguishable methyl protons, only nine distinct expectation values, $\langle\sigma^{m}_{\alpha}\rangle$ ($\alpha=x,y,z$, $m={\rm H_{1}},{\rm H_{2}},{\rm methyl}$), are accessible. Time-multiplexed readout is implemented through the standard NMR FID detection, which monitors the transverse magnetization via continuous weak measurement [45], requiring no extra experimental overhead [1]. The recorded signal can be written as
$S(t)\propto{\rm Tr}\left\{e^{t\mathcal{L}_{H}}\left[U\rho U^{\dagger}\right]O_{\rm FID}\right\},$ (4)
where $O_{\rm FID}=\sum_{m=5}^{9}\left(\sigma^{m}_{y}+i\sigma^{m}_{x}\right)$ is the collective transverse operator on protons. In the experiments, we apply a $\pi/2$ pulse before readout ($U=\prod_{n=5}^{9}R^{n}_{x}(\pi/2)$), and perform a Fourier transform on the FID signal to extract the reservoir response, selecting spectral peaks to form 653 readout features. Further details are provided in Supplemental Material [1].
Figure 3: High accuracy temporal prediction with NARMA benchmark. (a) Prediction performance (evaluated by $R^{2}$) on NARMA tasks with $n$ varying from 2 to 20, using quantum reservoirs with different configurations. Red and blue lines represent evolution durations of $\tau=$0.3\text{\,}\mathrm{s}$\$ and $0.01\text{\,}\mathrm{s}$, respectively, while solid and dashed lines indicate results obtained with time-multiplexed (TM) and single-time readout schemes. (b) Prediction results for NARMA tasks with $n=10$ and $20$, utilizing time-multiplexed readout with $\tau=$0.01\text{\,}\mathrm{s}$\$.
@@
where $\bar{y}=\frac{1}{L}\sum_{k=1}^{L}y_{k}$. For direct comparison with prior work, we also report the corresponding normalized mean square errors in Table 1; our best cases achieve error reductions of 1-2 orders of magnitude relative to recent experimental benchmarks [77, 86, 55]. Time-multiplexed readout markedly improves performance over single-time readout, especially for $\tau=$0.01\text{\,}\mathrm{s}$\$, where predictions are nearly perfect across all NARMA orders (see NARMA10 and 20 in Fig. 3(b) and additional results in Supplemental Material [1]). As the relaxation effects become more pronounced with longer evolution times, the information from high-order correlations is significantly diminished at $\tau=$0.3\text{\,}\mathrm{s}$\$, thereby reducing the benefit of time multiplexing.
In Supplemental Material [1], we show that the quantum reservoir markedly outperforms its classical spin-based counterpart, which yields an average $R^{2}<0.75$ across NARMA2 to NARMA20, highlighting the computational role of quantum correlations. We also provide a quantitative assessment of the protocol’s robustness against several experimental imperfections, offering guidance for implementing the scheme on various physical platforms.
@@
Realistic weather forecasting—To show the practical utility of our approach, we conduct weather forecasting on a daily climate dataset from Delhi, a benchmark commonly employed in the machine learning community [15]. Our analysis focuses on predicting the multivariate time series of temperature and humidity, ${\bf y}_{k}=\left[{\rm Temp}(k),{\rm Humi}(k)\right]$. We first perform single-step-ahead forecasting, where ${\bf y}_{k}$ serves as the $k$th input and ${\bf y}_{k+1}$ as the corresponding target. The dataset is divided into 374 washout steps, 600 training steps, and 600 testing steps. The normalized temperature and humidity data are encoded as angles $\theta_{k}$ and $\phi_{k}$ for global rotations applied to the proton and carbon nuclei, respectively, with an input interval of $\tau=$0.03\text{\,}\mathrm{s}$\$. Figure 4(a) presents the prediction results alongside the target values from the test set. Real weather data exhibit substantial stochastic fluctuations, which pose challenges for precise forecasting. Nevertheless, QRC predictions capture the overall trends and closely match local variations in several segments of the time series.